# From strange-quark tagging to fragmentation tagging with machine learning

### v2

@@ 1 Introduction
The second problem that we will address with the same tools is fragmentation tagging. By this we mean determining the fragmentation channel of a quark, e.g., identifying the type of $b$ hadron that was present in a $b$ jet. The motivation for developing fragmentation taggers comes from several directions. One is measurements of parton fragmentation functions (FFs), which describe how partons transform into hadrons metz2016parton ; albino2008akk . The FFs, which are determined by non-perturbative QCD matrix elements, describe the probability of a parton producing a specific hadron carrying a certain fraction of the parton momentum. Measuring FFs is useful for improving our understanding, or at least the description, of QCD dynamics, as well as for tuning Monte Carlo generators. While FF measurements can be done using specific clean decay channels of the corresponding hadrons (see, e.g., refs. seuster2006charm ; CMS-PAS-TOP-18-012 ), it is interesting to ask whether ML methods can help doing more inclusive measurements. Such measurements might be particularly useful at high $p_{T}$, where statistics is limited.
@@ 2.1 Event generation
We simulate dijet events in proton-proton collisions with a center-of-mass energy of $14$ TeV. We use MadGraph5 3.5.1 alwall2014automated to create the hard scattering process and Pythia 8.308 bierlich2022comprehensive to handle the parton showering, hadronization and decays. We define two kinematic regions for the jets (clustered as described below): $p_{T,{\rm jet}}>200$ GeV and $p_{T,{\rm jet}}>45$ GeV. These jet $p_{T}$ cuts follow generation-level cuts on the partons of $p_{T}>180$ GeV and $p_{T}>35$ GeV, respectively, and pseudorapidity $|\eta|<4$.
To obtain samples of $s$ and $d$ jets, we generate $s\bar{s}$ and $d\bar{d}$ events. We include only the $gg$ and $u\bar{u}$ initial states so that the only difference between the $s$ and $d$ jets in the resulting samples will be their flavor. We do it to ensure that the taggers rely solely on the intrinsic differences between the jets and not on differences between their $p_{T}$ and $\eta$ distributions in particular samples.
@@ 2.2 Detector simulation
We consider particles in the range $|\eta|<4$, which is approximately the range that will be covered by the ATLAS and CMS tracking detectors at the HL-LHC atlas2019expected ; cms2018expected . In addition, charged particles need to satisfy $p_{T}>0.5$ GeV. Particles are treated as stable if they do not decay within $1$ m from the beam axis and $3$ m along the beam axis from the interaction point. Based on them, we form the following detector-level objects with the help of the Monte Carlo truth information:
@@ 2.3 Reconstruction of $K_{S}$ and $\Lambda$ decays
The $K_{S}$ meson decays as $K_{S}\to\pi^{+}\pi^{-}$ with a 69% branching ratio, and $K_{S}\to\pi^{0}\pi^{0}$ with a 31% branching ratio. The $\Lambda$ baryon decays as $\Lambda\to p\pi^{-}$ with a 64% branching ratio, and $\Lambda\to n\pi^{0}$ with a 36% branching ratio. We attempt to reconstruct $K_{S}$ and $\Lambda$ as intermediate particles for decays to charged hadrons that occur at a distance greater than $0.5$ cm and less than $50$ cm from the beam axis. The lower bound helps to avoid confusion with prompt tracks that originate from the interaction point, while the upper bound represents the radius beyond which track reconstruction becomes essentially impossible due to an insufficient number of tracker layers that can be used. As mentioned above, we simulate the track reconstruction efficiency as a function of the track production radius based on ref. strebler2022expected . If both charged hadron tracks from the $K_{S}$ or $\Lambda$ decay are reconstructed, we remove them from the list of charged hadrons and store them as a reconstructed $K_{S}$ or $\Lambda$ object instead.
@@ 3 Basic discriminating variables
The final step in our analysis uses Receiver Operating Characteristic (ROC) curves. An ROC curve characterizes the discriminating power of applying a threshold to a given variable. It presents the signal efficiency ($\varepsilon_{s}$) vs. the background efficiency ($\varepsilon_{b}$) achieved at varying threshold values. In the present case, the signal efficiency corresponds to correctly identifying $s$-quark jets, and the background efficiency indicates the fraction of $d$-quark jets incorrectly identified as $s$-quark jets. For each of the variables, we construct an ROC curve and compute the Area Under the Curve (AUC). The AUC serves as a metric for the discriminative power of the variable. Random guessing would give an AUC of $0.5$, whereas an AUC of $1$ indicates perfect discrimination, and an AUC of 00 also signifies perfect discrimination but with an opposite threshold direction. Figure 4 presents the ROC curves for the ten most discriminative jet and constituent features. [^1] Features of constituents beyond the fifth most energetic one are not considered for this plot since they are less likely to be meaningful as individual variables and because they are not available in all jets. However, features of all the constituents of each jet will be made available to the advanced neural networks. The neutral hadron energy and photon energy, followed by the identity of the most energetic particle, whether it is a neutral hadron or a photon, show the highest/lowest AUC values, implying they are the most discriminative features. Figure 5 shows the distributions of the three most discriminative features identified in figure 4.
@@ 4.1 NN inputs
We represent each jet as a graph, implemented with Deep Graph Library (DGL) DBLP:journals/corr/abs-1909-01315 . The graph nodes represent the jet constituents. Each node’s features include the properties of both the jet and the constituent. We employ a fully-connected topology, where edges are formed between every pair of nodes. For an edge between node $i$ and node $j$, a vector is initialized with a list of the jet properties, the constituent properties from node $i$, and the constituent properties from node $j$. We leave it to each of the NNs to construct useful edge features based on these physical inputs. Reverse edges (between nodes $j$ and $i$), are included as well, with the order of the nodes in the vector swapped, to allow the independent flow of information in each direction. Each graph in our simulated dataset carries a label to denote the jet type: ‘0’ for $d$-quark jets or $b$-meson jets, and ‘1’ for $s$-quark jets or $b$-baryon jets.
@@ 4.3.1 Strange tagging
Figure 12 shows that all the NNs outperform the most discriminative individual features (cf. figure 4). However, all the models have very similar performance, which is also quite similar to what has been obtained with the simpler architectures explored in the past: BDTs nakai2020strange , CNNs nakai2020strange , LSTM RNNs erdmann2021maximum ; zeissner2021development and FNNs zeissner2021development . The sophisticated GAT and ParT architectures do not bring any improvement in performance, suggesting that the jet data does not contain much useful information beyond what can be captured by a simple combination of hand-crafted variables. This leads us to believe that achieving significantly better strange-tagging performance with the ATLAS and CMS detectors is unlikely, at least with the physical inputs that we assumed to be available and potentially relevant.
@@ 4.3.2 Fragmentation tagging
4.4 Robustness to measurement errors
@@ 5 Summary and discussion
In strange tagging, we found that the GAT and ParT architectures provided with the full jet constituent data did not perform significantly better than a simple MLP that combined the features of the jet and the leading constituent. This extends the result of ref. nakai2020strange , where it was shown that CNNs applied to jet images did not significantly outperform BDTs that used a small number of key whole-jet variables or even a single hand-crafted variable.
@@ A.2 Particle Transformer (ParT)
The scaled dot-product attention mechanism starts by transforming each element (token) of the input data through learned linear projections into three vectors: a query $Q$, a key $K$, and a value $V$. Conceptually, the query is like a ‘search term’ for identifying relevant information, the key acts as a ‘label’ for each data point, and the value contains the actual data that the mechanism ultimately focuses on based on the computed relevance. Attention scores are computed by calculating the dot product of the query of one token and the key of another, scaling it by $\sqrt{d_{k}}$, [^3] The scaling by $\sqrt{d_{k}}$ is done to prevent the dot product from becoming too large in magnitude (which would force the softmax function to give a number very close to 1), which can cause vanishing gradients during backpropagation vaswani2017attention . where $d_{k}$ is the dimensionality of each key and query, and applying the softmax function to obtain the weights that will eventually multiply the values:

### v3

@@ 1 Introduction
The second problem that we will address with the same tools is fragmentation tagging. By this we mean determining the fragmentation channel of a quark, e.g., identifying the type of $b$ hadron that was present in a $b$ jet. The motivation for developing fragmentation taggers comes from several directions. One is measurements of parton fragmentation functions (FFs), which describe how partons transform into hadrons metz2016parton ; albino2008akk . The FFs, which are determined by nonperturbative QCD matrix elements, describe the probability of a parton producing a specific hadron carrying a certain fraction of the parton momentum. Measuring FFs is useful for improving our understanding, or at least the description, of QCD dynamics, as well as for tuning Monte Carlo generators. While FF measurements can be done using specific clean decay channels of the corresponding hadrons (see, e.g., refs. seuster2006charm ; CMS-PAS-TOP-18-012 ), it is interesting to ask whether ML methods can help doing more inclusive measurements. Such measurements might be particularly useful at high $p_{T}$, where statistics is limited.
@@ 2.1 Event generation
We simulate dijet events in proton-proton collisions with a center-of-mass energy of $14$ TeV. We use MadGraph5 3.5.1 alwall2014automated to create the hard scattering process at the leading order and Pythia 8.308 bierlich2022comprehensive to handle the parton showering, hadronization and decays. We define two kinematic regions for the jets (clustered as described below): $p_{T,{\rm jet}}>200$ GeV and $p_{T,{\rm jet}}>45$ GeV. These jet $p_{T}$ cuts follow generation-level cuts on the partons of $p_{T}>180$ GeV and $p_{T}>35$ GeV, respectively, and pseudorapidity $|\eta|<4$.
To obtain samples of $s$ and $d$ jets, we generate $s\bar{s}$ and $d\bar{d}$ events. We include only the $gg$ and $u\bar{u}$ initial states so that the only difference between the $s$ and $d$ jets in the resulting samples will be their flavor. [^1] If we included the $d\bar{d}$ initial state, for example, its contribution would be different for the $d\bar{d}$ and $s\bar{s}$ final states in terms of both its size (relative to the $gg$ contribution) and kinematics, due to the presence of a $t$-channel diagram in the first case only. Including the $s\bar{s}$ initial state as well would not compensate for this effect because of the lower $s$-quark content in the proton. We do it to ensure that the taggers rely solely on the intrinsic differences between the jets and not on differences between their $p_{T}$ and $\eta$ distributions in particular samples.
@@ 2.2 Detector simulation
We consider particles in the range $|\eta|<4$, which is approximately the range that will be covered by the ATLAS and CMS tracking detectors at the HL-LHC atlas2019expected ; cms2018expected . In addition, charged particles need to satisfy $p_{T}>0.5$ GeV. Particles are treated as stable if they do not decay within $1$ m from the beam axis and $3$ m along the beam axis from the interaction point. [^2] This roughly describes the inner detector volume in ATLAS ATLAS:2008xda and CMS CMS:2008xjf . Beyond this volume, particles are absorbed in the calorimeters. Based on them, we form the following detector-level objects with the help of the Monte Carlo truth information:
@@ 2.3 Reconstruction of $K_{S}$ and $\Lambda$ decays
The $K_{S}$ meson decays as $K_{S}\to\pi^{+}\pi^{-}$ with a 69% branching ratio, and $K_{S}\to\pi^{0}\pi^{0}$ with a 31% branching ratio. The $\Lambda$ baryon decays as $\Lambda\to p\pi^{-}$ with a 64% branching ratio, and $\Lambda\to n\pi^{0}$ with a 36% branching ratio. We emulate the reconstruction of $K_{S}$ and $\Lambda$ as intermediate particles for decays to charged hadrons that occur at a distance greater than $0.5$ cm and less than $50$ cm from the beam axis. The lower bound helps to avoid confusion with prompt tracks that originate from the interaction point, while the upper bound represents the radius beyond which track reconstruction becomes essentially impossible due to an insufficient number of tracker layers that can be used. Within this volume, $K_{S}$ and $\Lambda$ decays can be easily identified using their highly displaced vertices and the possibility to reconstruct the invariant mass of the two products, as long as the tracks are reconstructed (see, e.g., refs. ATLAS:2023nze ; ATLAS:2024nbm ; CMS:2013zgf ). As mentioned above, we simulate the track reconstruction efficiency as a function of the track production radius based on ref. strebler2022expected . If both tracks from the $K_{S}$ or $\Lambda$ decay are reconstructed, we remove them from the list of charged hadrons and store them as a reconstructed $K_{S}$ or $\Lambda$ object instead.
@@ 3 Basic discriminating variables
To see to what extent the jet classification task can be accomplished by simply applying a cut on the distribution of any individual variable, we utilize Receiver Operating Characteristic (ROC) curves. An ROC curve characterizes the discriminating power of applying a threshold to a given variable. It presents the signal efficiency ($\varepsilon_{s}$) vs. the background efficiency ($\varepsilon_{b}$) achieved at varying threshold values. In the present case, the signal efficiency corresponds to correctly identifying $s$-quark jets, and the background efficiency indicates the fraction of $d$-quark jets incorrectly identified as $s$-quark jets. For each of the variables, we construct an ROC curve and compute the Area Under the Curve (AUC). The AUC serves as a metric for the discriminative power of the variable. Random guessing would give an AUC of $0.5$, whereas an AUC of $1$ indicates perfect discrimination, and an AUC of 00 also signifies perfect discrimination but with an opposite threshold direction. Figure 4 presents the ROC curves for the ten most discriminative jet and constituent features. [^3] Features of constituents beyond the fifth most energetic one are not considered for this plot since they are less likely to be meaningful as individual variables and because they are not available in all jets. However, features of all the constituents of each jet will be made available to the advanced neural networks. The neutral hadron energy and photon energy, followed by the identity of the most energetic particle, whether it is a neutral hadron or a photon, show the highest/lowest AUC values, implying they are the most discriminative features. Figure 5 shows the distributions of the three most discriminative features identified in figure 4.
@@ 4.1 NN inputs
We represent each jet as a graph, implemented with Deep Graph Library (DGL) DBLP:journals/corr/abs-1909-01315 . The graph nodes represent the jet constituents. Each node’s features include the properties of both the jet and the constituent. We employ a fully connected topology, where edges are formed between every pair of nodes. For an edge between node $i$ and node $j$, a vector is initialized with a list of the jet properties, the constituent properties from node $i$, and the constituent properties from node $j$. We leave it to each of the NNs to construct useful edge features based on these physical inputs. Reverse edges (between nodes $j$ and $i$), are included as well, with the order of the nodes in the vector swapped, to allow the independent flow of information in each direction. Each graph in our simulated dataset carries a label to denote the jet type: ‘0’ for $d$-quark jets or $b$-meson jets, and ‘1’ for $s$-quark jets or $b$-baryon jets.
@@ 4.3.1 Strange tagging
Figure 12 shows that all the NNs outperform the most discriminative individual features (cf. figure 4). However, all the models have very similar performance, which is also quite similar to what has been obtained with the simpler architectures explored in the past: BDTs nakai2020strange , CNNs nakai2020strange , LSTM RNNs erdmann2021maximum ; zeissner2021development and FNNs zeissner2021development . The sophisticated GAT and ParT architectures do not bring any improvement in performance, suggesting that the jet data does not contain much useful information beyond what can be captured by a simple combination of hand-crafted variables. This leads us to believe that achieving significantly better strange-tagging performance with the ATLAS and CMS detectors is unlikely, at least with the physical inputs that we assumed to be available and potentially relevant. We note that the significantly better performance expected at the FCC-ee (e.g., ref. Blekman:2024wyf ) relies to a large extent on charged hadron identification capabilities, which allow in particular distinguishing $K^{\pm}$ from $\pi^{\pm}$.
@@ 4.3.2 Fragmentation tagging
4.4 Robustness to measurement errors and parton shower modeling
@@
To evaluate sensitivity to parton shower modeling, we performed the following test. We trained the neural networks with the default Pythia parton shower model and then tested their performance on events generated with VINCIA showers Fischer:2016vfv . We found that this did not affect the classification performance noticeably, with changes in the AUC comparable to typical statistical variations of size $0.002$.
@@ 5 Summary and discussion
In strange tagging, we found that the GAT and ParT architectures provided with the full jet constituent data did not perform significantly better than a simple MLP that combined the features of the jet and the leading constituent. This extends the result of ref. nakai2020strange , where it was shown that CNNs applied to jet images did not significantly outperform BDTs that used a small number of key whole-jet variables or even a single hand-crafted variable. We summarize our results alongside analogous results from the literature in table 1.
Task Architecture Study AUC $1/\varepsilon_{b}$ for $\varepsilon_{s}=0.1$
strange tagging BDT Ref. nakai2020strange 0.63, 0.63 22, 17
CNN Ref. nakai2020strange 0.64, 0.64 24, 19
LSTM Refs. erdmann2021maximum ; zeissner2021development 0.60 14
MLP Ref. zeissner2021development 0.61 20
MLP This work 0.65, 0.65 23, 20
GAT This work 0.65, 0.65 24, 20
ParT This work 0.65, 0.65 23, 20
fragmentation tagging MLP This work 0.75, 0.72 95, 57
GAT This work 0.80, 0.78 140, 80
ParT This work 0.80, 0.78 150, 80
Table 1: Comparison of performance metrics. For ref. nakai2020strange , the two numbers given in each cell correspond to jets with $p_{T,{\rm jet}}\approx 45$ GeV and jets with $p_{T,{\rm jet}}>200$ GeV, respectively. Similarly, for our results, the two numbers correspond to jets with $p_{T,{\rm jet}}>45$ GeV and $p_{T,{\rm jet}}>200$ GeV, respectively. Small differences in performance between different papers should not be taken too seriously because of differences in sample selection and detector simulation.
@@ A.2 Particle Transformer (ParT)
The scaled dot-product attention mechanism starts by transforming each element (token) of the input data through learned linear projections into three vectors: a query $Q$, a key $K$, and a value $V$. Conceptually, the query is like a search term for identifying relevant information, the key acts as a label for each data point, and the value contains the actual data that the mechanism ultimately focuses on based on the computed relevance. Attention scores are computed by calculating the dot product of the query of one token and the key of another, scaling it by $\sqrt{d_{k}}$, [^5] The scaling by $\sqrt{d_{k}}$ is done to prevent the dot product from becoming too large in magnitude (which would force the softmax function to give a number very close to 1), which can cause vanishing gradients during backpropagation vaswani2017attention . where $d_{k}$ is the dimensionality of each key and query, and applying the softmax function to obtain the weights that will eventually multiply the values: